OpenStudy (anonymous):

|{ -x+y+2z=7 <|{ 2x-3y-2z=-4 |{ x -z=-2 possible answers are (3,3,5) (3,0,5) (-3,0,5) (3,0,-5)

OpenStudy (anonymous):

the last one is the answer????

OpenStudy (anonymous):

or the one before the last

OpenStudy (anonymous):

ok lets see.

OpenStudy (anonymous):

last one doesn't work for lat equation. because 3-(-5) = 3+5 = 8

OpenStudy (anonymous):

third one doesn't work for last equation because -3-5=-8

OpenStudy (anonymous):

must be one of the first two. lets find it

OpenStudy (anonymous):

-3+3+10=7 <|{ 6-9-10=-4 |{ 3 -5=-2

OpenStudy (anonymous):

thats not it

OpenStudy (anonymous):

nope it is not first one

OpenStudy (anonymous):

-3+0+10=7 6-0-10=-4 3 5=-2

OpenStudy (anonymous):

must be (3,0,5)

OpenStudy (anonymous):

thats not it

OpenStudy (anonymous):

yes, late equation is x - z = -2 so 3-5=-2 is correct

OpenStudy (anonymous):

-3+0+10=7 6-0-10=-4 -3 5=-2

OpenStudy (anonymous):

-3+7=10 6-10=-4 3-5=-2

OpenStudy (anonymous):

the first one hast to be 7

OpenStudy (anonymous):

last equation is x - z = -2 you wrote -x+z

OpenStudy (anonymous):

no -5 cancelled out

OpenStudy (anonymous):

yes it is if you use x=3,y=0,z=5

OpenStudy (anonymous):

-3+0+2*5=-3+10=7

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

second equation is 2x-3y-2z = -4 we get 2*3-0-2*5=6-10=-4

OpenStudy (anonymous):

and third is x-z=-2 we get 3-5=-2

OpenStudy (anonymous):

so they all work for (3,0,5)

OpenStudy (anonymous):

your right i am going to choose yours

OpenStudy (anonymous):

good because it is true!